# A model of thermophoresis of colloidal proteins in water using non-Fickian diffusion currents

### v1

@@ IV Discussion
On the other hand, we have also taken into consideration an explicit temperature dependence of diffusivity via the temperature dependence of the dynamic viscosity $\eta(T)$. Here, keeping in mind that this temperature dependence of the deviation of the dynamic viscosity from that of pure water can occur based on the interaction of the particle’s surface with the nearby liquid layers, we have considered this dependence to be proportional to the local thermal expansion coefficient of water. We had to fix it this way, considering that if any solvent observable couples to the activity on the particle surface, that should, in general, be the density of the solvent. And these considerations worked quite well in matching the empirical profile of $S^{\prime}_{T}$ for the test cases we have considered. We have considered these test cases, recognising that, within the scope of the present model, strong electrostatic interactions cannot be captured. But, what we want to achieve at present is to produce observed profiles of $S^{\prime}_{T}$ using not more than three parameters, as the same number of parameters is considered in the empirical relation eq.(16).
@@
    The coupling between the solvation and diffusion through the correction to dynamic viscosity, as seen by the colloidal particle, is a crucial ingredient in our model. Interestingly, the parameter $c$ has a negative sign for BLGA and Poly-L-Lysine, but not for Lysozyme. The positive $c$ for Lysizyme may indicate the occurrence of slips in the solvation layer, whereas the negative sign of the $c$ for BLGA and Poly-L-Lysine could point to a reduction of mobility due to stronger bonding.
@@
    The hydrodynamic effects bringing in the other routes of temperature dependence of the diffusivity via its concentration dependence can flip the sign of the $S_{T}$ depending upon $\beta>1$ or $\beta<1$. In our test cases, with colloidal proteins $\beta\simeq 1.3-1.38$. Apart from changing the sign of $S_{T}$, the $|1-\beta|$ sets the scale for the Soret coefficient. Theoretical results on the distance $r$ dependence of diffusivity at an interface, based on hydrodynamic considerations, apply to simple geometries where setting the boundary conditions is possible [58, 59, 60]. Considering $C\sim r^{-3}$, in the present case, roughly considering the $\beta\sim 4/3$ indicates an $(r/r_{0})^{-4}$ dependence of the increase in viscosity on the average distance of the interface from the surface of a particle. Here, $r_{0}$ is the corresponding length scale to $C_{0}$.
@@
    It is observed that the effect of the solvation force in our model is relatively innocuous as compared to the diffusion terms. Within the scope of our model, where the solvation force is proportional to $T(dT/dx)$, based on the equation of state considered, it produces a vertical translation to the Soret coefficient curve. Therefore, to model solvation forces in more complex scenarios, one may look for newer forms of the local equation of state. For this purpose, our model sets a benchmark.
@@
    The power of the physical considerations underlying our present model is also evident from Fig.5(c), where we show practically an exact match to $C_{sat}(T)$ experimentally obtained. We have presented this plot to underscore that our consideration of Chapman’s non-Fickian diffusion current in the presence of a temperature gradient effectively captures the local equilibrium distribution at a particular temperature. This underscores the fact that the Chapman-Itô form of the current density is the generalised form for inhomogeneous diffusion.

### v2

@@ IV Discussion
On the other hand, we have also taken into consideration an explicit temperature dependence of diffusivity via the temperature dependence of the dynamic viscosity $\eta(T)$. Here, keeping in mind that this temperature dependence of the deviation of the dynamic viscosity from that of pure water can occur based on the interaction of the particle’s surface with the nearby liquid layers, we have considered this dependence to be proportional to the local thermal expansion coefficient of water. We had to fix it this way, considering that if any solvent observable couples to the activity on the particle surface, that should, in general, be the density of the solvent. And these considerations worked quite well at matching the empirical profile of $S^{\prime}_{T}$ and the experimental data for the test cases we have considered. We have considered these test cases, recognising that, within the scope of the present model, strong electrostatic interactions cannot be captured. But, what we want to achieve at present is to produce observed profiles of $S^{\prime}_{T}$ using not more than three parameters, as the same number of parameters is considered in the empirical relation eq.(16).
It is important to note that there may be very different parameter regimes in $a,c,\beta$ in which a similar curve-fitting could be obtained for $S_{T}$. However, only the current parameter range used to fit the curves makes physical sense for the following reasons. (a) Since the parameter $a$ provides a vertical shift to the $S_{T}$ curve according to our model, it cannot be of any other order than OO($10^{-2}$) for the test cases. (b) The magnitude of the parameter $c$ should be less than unity in order to avoid having the correction to the $\eta_{w}(T)$ of the same order as it. (c) Since the parameter $\beta$ flips the $S_{T}$ curve, it is $10>\beta>1$ to keep the order of magnitude of $S_{T}$ in the experimental regime besides getting the flip. We have checked that, corresponding to a relatively higher value of the $\beta$, there are no values of the parameters $a$ and $c$ in the permissible limit, and, moreover, the $C^{*}_{sat}(T)$ for Lysozyme is getting a poor fit.
@@
    The coupling between the solvation and diffusion through the correction to dynamic viscosity, as seen by the colloidal particle, is a crucial ingredient in our model. The positive $c$ value across all test cases indicates the presence of slips in the solvation layer, which is likely due to changes in hydrogen bonding.
@@
    The hydrodynamic effects bringing in the other routes of temperature dependence of the diffusivity via its concentration dependence can flip the sign of the $S_{T}$ depending upon $\beta>1$ or $\beta<1$. In our test cases, with colloidal proteins $\beta\simeq 1.179-1.265$. Apart from changing the sign of $S_{T}$, the $|1-\beta|$ sets the scale for the Soret coefficient. Theoretical results on the distance $r$ dependence of diffusivity at an interface, based on hydrodynamic considerations, apply to simple geometries where setting the boundary conditions is possible [67, 68, 69]. Considering $C\sim r^{-3}$, in the present case, roughly considering the $\beta\sim 5/4$ indicates an $(r/r_{0})^{-3.75}$ dependence of the increase in viscosity on the average distance of the interface from the surface of a particle. Here, $r_{0}$ is the corresponding length scale to $C_{0}$.
@@
    It is observed that the solvation current and the non-Fickian current are the dominant contributors to the density instability in the presence of the temperature gradient for the test cases considered. Within the scope of our model, where the solvation force is proportional to $T(dT/dx)$, based on the equation of state considered, it produces a vertical translation to the Soret coefficient curve. It needs to be seen whether the same local equation-of-state structure can work for other, more complex cases.
@@
    The power of the physical considerations underlying our present model is also evident from Fig. 5(c), where we show practically an exact match to $C^{*}_{sat}(T)$ experimentally obtained. We have presented this plot to underscore that our consideration of Chapman’s non-Fickian diffusion current in the presence of a temperature gradient effectively captures the local equilibrium distribution at a particular temperature. This underscores the fact that the Chapman-Itô form of the current density is the generalised form for inhomogeneous diffusion.
@@ Acknowledgements.
Appendix
The model parameters $(a,\beta,c)$ are determined by fitting the theoretical (model) prediction of the Soret coefficient to the experimental data using least square minimisation. For a given set of parameters, the model yields a temperature-dependent Soret coefficient $S_{T}^{\text{model}}(T;a,\beta,c)$, which is compared against the experimentally measured values $S_{T}^{\text{exp}}(T)$.
We define a least-squares error function
$\mathcal{E}(a,\beta,c)=\sum_{i}\left[S_{T}^{\text{model}}(T_{i};a,\beta,c)-S_{T}^{\text{exp}}(T_{i})\right]^{2},$
where the sum runs over all experimental temperature points $T_{i}$. The optimal parameter set $(a^{*},\beta^{*},c^{*})$ is obtained by minimizing $\mathcal{E}$ are displayed in table 1 in main text, where they are denoted by $(a,\beta,c)$ respectively.
To explore the structure of the parameter space and assess parameter sensitivity, we display the error landscape below by showing pairs of parameters while keeping the third parameter fixed at its optimal value. A contour plot of the logarithm of the error function, $\log\mathcal{E}$, is used for visualisation purposes.
Refer to caption(a)(b)(c)
Figure A1: Contour plot of $\log\mathcal{E}$ showing the error landscape in parameter space for Lysozyme protein.
Refer to caption(a)(b)(c)
Figure A2: Contour plot of $\log\mathcal{E}$ showing the error landscape in parameter space for BLGA protein.
Refer to caption(a)(b)(c)
Figure A3: Contour plot of $\log\mathcal{E}$ showing the error landscape in parameter space for Poly-L-Lysine protein.